OpenStudy (anonymous):

Carpet stain protector costs $0.65 per square yard to apply. How much will it cost to apply the protectant to a 16-foot × 18-foot carpet? A. $187.20 B. $62.40 C. $20.80 D. $96.00 please show me how to solve it by steps

OpenStudy (anonymous):

18' X 16' = 288sq ft 288 sq ft / 9 = 32 sq yds 32 sq yds X $0.65 = $20.80

OpenStudy (anonymous):

oh so thats how easy it is... wait you changed it XD @MagicallyFarting .. so its really $20.80? not 187 :)

OpenStudy (anonymous):

1 foot = 1/3 yard

OpenStudy (anonymous):

Yeah haha xD Sorry I derped the first moment.

OpenStudy (anonymous):

alrighty thanks guys for the help :D @MagicallyFarting & @ganymede

OpenStudy (anonymous):

oh wait @MagicallyFarting where did the 9 came from?

OpenStudy (anonymous):

1 foot squared = 1/9 yards

OpenStudy (anonymous):

so for 288 ft sq/ you just have to plug in 9?

OpenStudy (anonymous):

If it were 288 feet, knowing 1 foot = 1/3 of a yard, for every foot your will have 1/3 of a yard. So, if you had 288 feet your would have 1/3 as many yards. So 288 * 1/3 = whatever.

OpenStudy (anonymous):

But we are dealing with feet squared. 1 foot squared = 1/9 yards squared. So if you had 288 feet squared you would have 1/9 as many yards squared. So 288 * 1/9 = 32 yards squared.

OpenStudy (anonymous):

288 * 1/9 is the same as 288/9. I hope that helps!

OpenStudy (anonymous):

thanks now i understand thank you